OpenStudy (anonymous):

Find the limit. Use l'Hospital's Rule if appropriate lim x->0 (x(6^x))/((6^x)-1)

OpenStudy (anonymous):

not that easy to read

OpenStudy (anonymous):

\[\frac{6x^x}{6^x-1}\]?

OpenStudy (anonymous):

the bottom is right the top is x6^x

OpenStudy (anonymous):

\[\huge \lim_{x\to 0}\frac{x6^x}{6^x-1}\]

OpenStudy (anonymous):

yeah thats it

OpenStudy (anonymous):

ok it is in the for \(\frac{0}{0}\) take the derivative top and bottom separately

OpenStudy (anonymous):

*form

OpenStudy (anonymous):

you know the derivative of these?

OpenStudy (anonymous):

give me one second to figure them out

OpenStudy (anonymous):

here is a huge hint if you need it for the derivative of the numerator you need the product rule also \[\frac{d}{dx}b^x=b^x\ln(b)\]

OpenStudy (anonymous):

ln(6)x6^x+6^x/ln(6)6^x is that right?

OpenStudy (anonymous):

looks right

OpenStudy (anonymous):

now replace \(x\) by \(0\) and you are done

OpenStudy (anonymous):

i get \[\frac{1}{\ln(6)}\] but i didn't do it with paper, just looked see if that is right

OpenStudy (anonymous):

thats what i got thanks for the help

OpenStudy (anonymous):

yw